# ViM-VQ: Efficient Post-Training Vector Quantization for Visual Mamba

### v1

@@ 4.2 Fast Convex Combination Optimization
$\mathcal{C}_{o,i/d}=\{c_{o,i/d}^{1},\dots,c_{o,i/d}^{n}\}=\arg\min_{k}^{n}\|w_{o,i/d}-c(k)\|_{2}^{2},$ (5)
where $\mathcal{C}_{o,i/d}$ is candidate codewords of sub-vector $w_{o,i/d}$. Next, we compute soft assignment ratios $\mathcal{R}{o,i/d}$ for these candidate codewords:
$\mathcal{R}_{o,i/d}=\{r_{o,i/d}^{1},\dots,r_{o,i/d}^{n}\}=\{\dots,\frac{\exp\left({z_{o,i/d}^{n}}\right)}{\sum_{j=1}^{n}\exp\left({z_{o,i/d}^{j}}\right)}\},$ (6)
subject to the softmax constraint $\sum_{n}r_{o,i/d}^{n}=1$. Here, $z_{o,i/d}^{n}$ represents the unnormalized score for each ratio. In this way, the learnable $\mathcal{C}_{o,i/d}$ along with the learnable $\mathcal{R}_{o,i/d}$ form differentiable convex combinations that produces the quantized weight $\widehat{\mathbf{W}}$ as follows:
$\widehat{\mathbf{W}}\!=\!\\ \begin{bmatrix}\widehat{w}_{1,1}\!=\!\sum\mathcal{C}_{1,1}\mathcal{R}_{1,1}&\!\cdots\!&\widehat{w}_{1,i/d}\!=\!\sum\mathcal{C}_{1,i/d}\mathcal{R}_{1,i/d}\\ \widehat{w}_{2,1}\!=\!\sum\mathcal{C}_{2,1}\mathcal{R}_{2,1}&\!\cdots\!&\widehat{w}_{2,i/d}\!=\!\sum\mathcal{C}_{2,i/d}\mathcal{R}_{2,i/d}\\ \vdots&\ddots&\vdots\\ \widehat{w}_{o,1}\!=\!\sum\mathcal{C}_{o,1}\mathcal{R}_{o,1}&\!\cdots\!&\widehat{w}_{o,i/d}\!=\!\sum\mathcal{C}_{o,i/d}\mathcal{R}_{o,i/d}\\ \end{bmatrix}.$ (7)
For the initialization of each $\mathcal{R}_{o,i/d}$, we optimize $z_{o,i/d}^{n}$ via gradient descent to minimize the reconstruction error between $\widehat{\mathbf{W}}$ and $\mathbf{W}$:
$z_{o,i/d}^{n}=z_{o,i/d}^{n}-\eta\frac{\partial\mathcal{L_{\text{init}}}}{\partial z_{o,i/d}^{n}},\quad\mathcal{L_{\text{init}}}=\|\mathbf{W}-\widehat{\mathbf{W}}\|_{2}^{2},$ (8)
where $\eta$ denotes the learning rate. This initialization ensures that $\widehat{\mathbf{W}}$ is initially close to $\mathbf{W}$, thus preventing severe gradient oscillations during subsequent calibration.
Therefore, the convex combination optimization facilitates the quantized sub-vectors’ search for the optimal codeword. How can we ensure that optimal codewords are contained within the convex hulls? The quantized sub-vectors are often distant from those irrelevant candidate codewords with near-zero ratios, which indicates that the optimal codeword is not located near these irrelevant candidates. Consequently, we can replace them with new candidate codewords that are closest to the quantized sub-vector:
$c_{o,i/d}^{n}|_{r_{o,i/d}^{n}\approx 0}=\arg\min_{k}\|\widehat{w}_{o,i/d}-c(k)\|_{2}^{2}\notin\mathcal{C}_{o,i/d}$ (9)
where $r_{o,i/d}^{n}\approx 0$ is smaller than $\lambda$ (e.g., 1e-2). This adaptive replacement procedure dynamically shifts the convex hulls towards more optimal codewords without introducing significant perturbations to the network’s outputs.
@@
  • Unlike global weighted average methods [1, 42], which incur high calibration costs and memory overhead, or local weighted average methods [4], which yield suboptimal quantization, our approach efficiently balances these aspects. Our algorithm utilizes fewer codewords and soft assignment ratios to reduce GPU memory usage and calibration steps. Simultaneous updates of convex combinations and convex hulls accelerate the search for optimal codewords, achieving superior performance.
@@
  • Our algorithm is a differentiable process and does not require additional computation such as Hessian trace [6, 3, 26, 8] or SVD [18, 38, 45].
@@ 4.3 Incremental Vector Quantization
Inspired by INQ [43], we propose an incremental vector quantization strategy. As illustrated in Figure 4, our strategy confirms candidate codewords with soft assignment ratios close to 1.0 (e.g. 0.99) as optimal codewords at each calibration step:
$\widehat{w}_{o,i/d}=c_{o,i/d}^{n}|_{r_{o,i/d}^{n}\approx 1}.$ (11)
Compared to previous methods that perform a one-time assignment conversion, our incremental strategy significantly reduces truncation errors at each calibration step. This allows the weights that remain in calibration to effectively recover the network’s accuracy. Moreover, our method ensures consistent network performance between the calibration and inference phases.
The incremental vector quantization strategy introduces three objective functions to efficiently identify and confirm optimal codewords. Specifically, given the input $\mathbf{x}$ and the target $\mathbf{y}$ from the calibration dataset for the quantized ViMs $\mathbf{\epsilon}_{q}$, the task objective function is defined as:
@@
where $b_{fp}^{l}$ and $b_{q}^{l}$ represent the features of $l$-th block from $\mathbf{\epsilon}_{fp}$ and $\mathbf{\epsilon}_{q}$, respectively.
Additionally, we introduce a regularization term $\mathcal{L}_{r}$ to encourage rapid convergence of the differentiable convex combinations:
$\mathcal{L}_{r}=\left(\frac{d}{o\times i}\right)\times\sum_{o,i/d}\left(\mathcal{R}_{o,i/d}\times\left(1-\mathcal{R}_{o,i/d}\right)\right).$ (14)
The final combined objective function $\mathcal{L}$ is formulated as:
@@
where $\mathcal{L}_{r}$ is selectively included only when it increases compared to the previous iteration. This selective incorporation ensures stable updates and convergence of the differentiable convex combinations, ultimately accelerating the search for optimal codewords.
@@ 5 Experiment
Our experiments are conducted on two Visual Mamba networks (ViMs) [24]: Vision Mamba (Vim) [44] and VMamba [25]. The experiments cover four main computer vision tasks: image classification, semantic segmentation, object detection, and instance segmentation. Our primary goal is to investigate the trade-off between model weight quantization and accuracy. Therefore, activation quantization is disabled in all experiments. We sample 100 images from each training category to form the calibration set, using a batch size of 128. All experiments are conducted on a single NVIDIA A6000 GPU with 48GB of memory.
Hyperparameters: The codebook sizes for each layer in the 3/2/1-bit quantization are set to $2^{6}\times 2$, $2^{8}\times 4$, and $2^{8}\times 8$, respectively. Soft assignments ratios are updated by the Adamax optimizer with a fixed learning rate of $5\times 10^{-2}$ without additional hyperparameter tuning. Dataset processing and other hyperparameters are identical to those of the original ViMs. Consistent with all baselines, we quantize only the linear projection layers within the ViMs blocks, as these layers contribute significantly to memory usage.
@@
Model Method Top-1 Accuracy (%)
@@ 5.1 Image Classification
Table 1 presents the image classification performance on three different-scale Vision Mamba (Vim-T/S/B) evaluated on ImageNet-1K. Uniform quantization (UQ) methods collapse in 2-bit quantization due to significant quantization errors, even with the latest Vim-specific approaches, such as Mambaquant and PTQ4VM. Across all three networks and various bit-widths, our ViM-VQ consistently achieves state-of-the-art quantization performance, maintaining the highest Top-1 accuracy. Specifically, for ViM-T, ViM-VQ achieves accuracies of 74.79% (3-bit), 72.17% (2-bit), and 69.93% (1-bit), clearly outperforming other advanced VQ methods like DKM and VQ4DiT. Similar trends appear in ViM-S and ViM-B, where ViM-VQ maintains performance very close to the full-precision baseline, demonstrating robustness even under extremely low-bit quantization.
@@ 5.2 Semantic Segmentation
Table 2 presents the semantic segmentation results of VMamba-T on the ADE20K dataset using UperNet as the segmentation framework. At 2-bit quantization, ViM-VQ achieves an mIoU of 44.4% (single-scale) and 45.4% (multi-scale), attaining performance comparable to full-precision Swin-T and ConvNeXt-T. Even under aggressive 1-bit quantization, ViM-VQ maintains an mIoU above 40%, demonstrating its robustness in extremely low-bit settings.
@@ 5.3 Object Detection and Instance Segmentation
Table 3 presents the object detection and instance segmentation performance on MSCOCO using Mask R-CNN as the detection framework. At 2-bit quantization, ViM-VQ consistently outperforms full-precision Swin-T and ConvNeXt-T, achieving 46.0% box AP and 41.8% mask AP. Even under the extreme condition of 1-bit quantization, ViM-VQ significantly mitigates performance degradation, obtaining 40.9% box AP and 35.8% mask AP.
@@
Table 5: Calibration efficiency of 2-bit quantization of tiny-scale Vision Mamba (Vim-T) on ImageNet-1K. $\mathbf{C}$, $n$, $E$, ”Mem” represent the codebook size, the number of candidate codewords, the number of epochs, and the GPU memory used for calibration, respectively. ”Acc-C” and ”Acc-I” represent the Top-1 accuracy (%) during calibration and inference.
@@ 5.4 Ablation Study
Table 4 presents an ablation study on the contribution of each component in our proposed ViM-VQ framework. The basic VQ method suffers from substantial quantization errors due to weight outliers, leading to a significant accuracy drop to 33.18%. By introducing the differential convex combination optimization algorithm, the accuracy improves by 38.24%, surpassing the performance of both DKM and VQ4DiT. This result highlights that updating both the convex combinations and the convex hulls effectively facilitates the search for the optimal codewords. Furthermore, incorporating the incremental vector quantization strategy provides an additional 0.75% improvement, demonstrating that the truncation error induced during the confirmation of optimal codewords is effectively mitigated.
@@
We develop a customized CUDA kernel for efficient matrix–vector multiplication. In our implementation, the codebooks are stored in constant GPU memory, while the assignments are fully loaded into CUDA blocks, enabling each thread to decode its corresponding assignment in parallel. The input is sequentially multiplied by the decoded weights, with the results accumulated in registers. As shown in Table 6, our CUDA kernel reduces inference time from 357s to 328s, bringing the speed of the quantized model similar to that of the original ViMs.
@@ 6 Conclusion
In this work, we analyze key challenges encountered when applying vector quantization (VQ) to Visual Mamba networks (ViMs). First, significant weight outliers in ViMs lead to severe quantization errors. Second, existing VQ methods applied to ViMs require substantial calibration steps, high memory consumption, and yield suboptimal performance. Additionally, truncation errors further degrade inference performance. We propose an efficient post-training VQ method tailored for ViMs, termed ViM-VQ. ViM-VQ incorporates a fast convex combination optimization algorithm for optimal codeword search and an incremental vector quantization strategy to confirm optimal codewords while mitigating truncation errors. Experimental results demonstrate that ViM-VQ significantly outperforms existing approaches across diverse visual tasks, offering a practical solution for deploying ViMs on edge devices.

### v2

@@ 4.2 Fast Convex Combination Optimization
$\mathcal{C}_{o,i/d}\kern-2.0pt=\kern-2.0pt\left\{c_{o,i/d}^{1},\dots,c_{o,i/d}^{n}\right\}\kern-2.0pt=\kern-2.0pt\operatorname{Top}_{n}(-\|w_{o,i/d}-c(k)\|_{2}^{2}),$ (4)
where $\mathcal{C}_{o,i/d}$ is the set of candidate codewords for sub-vector $w_{o,i/d}$. Next, we compute the soft assignment ratios $\mathcal{R}_{o,i/d}\kern-2.0pt=\kern-2.0pt\left\{r_{o,i/d}^{1},\dots,r_{o,i/d}^{n}\right\}$ for these candidates using the softmax function on a set of learnable scores $\{z_{o,i/d}^{m}\}_{m=1}^{n}:$
$r_{o,i/d}^{m}=\frac{\exp\left({z_{o,i/d}^{m}}\right)}{\sum_{j=1}^{n}\exp\left({z_{o,i/d}^{j}}\right)},\quad\text{where}\sum_{m=1}^{n}r_{o,i/d}^{m}=1.$ (5)
Here, $z_{o,i/d}^{m}$ represents the unnormalized score for the $m$-th ratio. In this way, the learnable $\mathcal{C}_{o,i/d}$ along with the learnable $\mathcal{R}_{o,i/d}$ form differentiable convex combinations that produce the quantized weight $\widehat{\mathbf{W}}$ as follows:
$\widehat{\mathbf{W}}\kern-5.0pt=\kern-5.0pt\begin{bmatrix}\widehat{w}_{1,1}\!=\!\sum\mathcal{C}_{1,1}\mathcal{R}_{1,1}&\!\cdots\!&\widehat{w}_{1,i/d}\!=\!\sum\mathcal{C}_{1,i/d}\mathcal{R}_{1,i/d}\\ \widehat{w}_{2,1}\!=\!\sum\mathcal{C}_{2,1}\mathcal{R}_{2,1}&\!\cdots\!&\widehat{w}_{2,i/d}\!=\!\sum\mathcal{C}_{2,i/d}\mathcal{R}_{2,i/d}\\ \vdots&\ddots&\vdots\\ \widehat{w}_{o,1}\!=\!\sum\mathcal{C}_{o,1}\mathcal{R}_{o,1}&\!\cdots\!&\widehat{w}_{o,i/d}\!=\!\sum\mathcal{C}_{o,i/d}\mathcal{R}_{o,i/d}\\ \end{bmatrix}\kern-3.0pt.$ (6)
The combinations can be updated via gradient descent:
$\mathcal{C}_{o,i/d}$ $\leftarrow\mathcal{C}_{o,i/d}-O(\nabla_{\mathcal{C}_{o,i/d}}\mathcal{L_{\text{init}}},\theta_{c}),$ (7)
$\mathcal{R}_{o,i/d}$ $\leftarrow\mathcal{R}_{o,i/d}-O(\nabla_{\mathcal{R}_{o,i/d}}\mathcal{L_{\text{init}}},\theta_{r}),$
where $O(,)$ is the optimizer with hyperparameter $\theta$. For the initialization of them, we optimize their corresponding codewords $c_{o,i/d}^{m}$ and ratios $r_{o,i/d}^{m}$ to minimize the reconstruction error $\mathcal{L_{\text{init}}}=\|\mathbf{W}-\widehat{\mathbf{W}}\|_{2}^{2}$. This initialization ensures that $\widehat{\mathbf{W}}$ is initially close to $\mathbf{W}$, thus preventing severe gradient oscillations during subsequent calibration.
The convex combination optimization facilitates the search for optimal codewords. To ensure optimal codewords are contained within effective convex hulls, we introduce an adaptive replacement strategy. The key insight is that if a current quantized sub-vector $\widehat{w}_{o,i/d}$ is distant from certain candidates (i.e., those with near-zero ratios), the optimal codeword is likely not near them. Consequently, we replace these irrelevant candidates with new candidate codewords that are closest to $\widehat{w}_{o,i/d}$:
$c_{o,i/d}^{m}\leftarrow\underset{c(k)\in\mathbf{C}\setminus\mathcal{C}_{o,i/d}}{\operatorname{argmin}}\|\widehat{w}_{o,i/d}-c(k)\|_{2}^{2},$ (8)
where $r_{o,i/d}^{m}\approx 0$ is smaller than $\lambda$ (e.g., 1e-2). This adaptive replacement procedure dynamically shifts the convex hulls towards more optimal codewords without introducing significant perturbations to the network’s outputs.
@@
  • Unlike global methods [1, 42], which incur high calibration costs and memory overhead, or local methods [4], which yield suboptimal quantization, our approach strikes an effective balance. The simultaneous optimization of convex combinations (ratios) and convex hulls (candidate codewords) accelerates the search process and reduces GPU memory usage, resulting in superior performance.
@@
  • Our algorithm is a fully differentiable process and does not require expensive secondary computations such as Hessian trace [6, 3, 26, 8] or SVD [18, 38, 45].
Figure 4: Incremental vector quantization strategy of ViM-VQ. High-ratio soft assignments are converted to hard assignments to confirm optimal codewords, while remaining soft assignments compensate for truncation errors during calibration.
@@ 4.3 Incremental Vector Quantization
Inspired by INQ [43], we propose an incremental vector quantization strategy. As illustrated in Figure 4, our strategy confirms candidate codewords with soft assignment ratios close to 1.0 (e.g., $>$ 0.99) as optimal codewords at each calibration step:
$\widehat{w}_{o,i/d}\leftarrow c_{o,i/d}^{m},\quad a_{o,i/d}\leftarrow\text{index}(c_{o,i/d}^{m}).$ (10)
Compared to previous methods that perform a one-time assignment conversion, our incremental strategy significantly reduces truncation errors at each calibration step. This allows the remaining weights still undergoing calibration to compensate for these quantization errors, thereby effectively recovering the network’s accuracy. Moreover, our method ensures consistent network performance between the calibration and inference phases.
The strategy employs a combination of three objective functions to efficiently identify and confirm optimal codewords. Specifically, given the input $\mathbf{x}$ and the target $\mathbf{y}$ from the calibration dataset for the quantized ViM model $\mathbf{\epsilon}_{q}$, the task objective function is defined as:
@@
where $b_{fp}^{l}$ and $b_{q}^{l}$ represent the output features of the $l$-th block from the full-precision model $\mathbf{\epsilon}_{fp}$ and the quantized model $\mathbf{\epsilon}_{q}$, respectively.
Additionally, we introduce a regularization term $\mathcal{L}_{r}$ to encourage rapid convergence of convex combinations:
$\mathcal{L}_{r}=\left(\frac{d}{o\times i}\right)\sum_{o,i/d}\sum_{m=1}^{n}r_{o,i/d}^{m}\left(1-r_{o,i/d}^{m}\right).$ (13)
The final objective function $\mathcal{L}$ is formulated as:
@@
where $\mathcal{L}_{r}$ is selectively included only when it increases compared to the previous iteration. This selective incorporation ensures stable updates and convergence of the differentiable convex combinations.
@@ 5 Experiment
Our experiments are conducted on two Visual Mamba networks (ViMs) [24]: Vision Mamba (Vim) [44] and VMamba [25]. The experiments cover four main computer vision tasks: image classification, semantic segmentation, object detection, and instance segmentation. Our primary goal is to investigate the trade-off between model weight quantization and accuracy. Therefore, activation quantization is disabled in all experiments. We sample 100 images per category from the training set to form the calibration set, using a batch size of 128. All experiments are conducted on a single NVIDIA A6000 GPU with 48GB of memory.
Hyperparameters: The codebook sizes for each layer in the 3/2/1-bit quantization are set to $2^{6}\times 2$, $2^{8}\times 4$, and $2^{8}\times 8$, respectively. The codebooks and ratios of soft assignments are updated using the Adamax optimizer with a fixed learning rate of $1\times 10^{-5}$ and $5\times 10^{-2}$, respectively, without additional hyperparameter tuning. Dataset processing and other hyperparameters follow the original settings of ViMs. Consistent with all baselines, we quantize only the linear projection layers within the ViMs blocks, as they contribute significantly to memory usage.
@@
Model (Params) Method Top-1 Accuracy (%)
@@ 5.1 Image Classification
Table 1 presents the image classification performance on three Vision Mamba models of different scales (Vim-T/S/B) evaluated on ImageNet-1K. Uniform quantization (UQ) methods collapse at 2-bit quantization due to significant quantization errors, even with the latest Vim-specific approaches, such as MambaQuant and PTQ4VM. Across all model scales and bit-widths, our ViM-VQ consistently achieves state-of-the-art quantization performance, maintaining the highest Top-1 accuracy. Specifically, for ViM-T, ViM-VQ achieves accuracies of 74.79% (3-bit), 72.17% (2-bit), and 69.93% (1-bit), clearly outperforming other advanced VQ methods like DKM and VQ4DiT. These trends hold for the larger ViM-S and ViM-B models, where ViM-VQ’s performance remains remarkably close to the full-precision baseline, demonstrating its robustness even under extremely low-bit conditions.
@@ 5.2 Semantic Segmentation
As shown in Table 2, our method demonstrates strong performance on the ADE20K semantic segmentation benchmark. At 2-bit quantization, ViM-VQ achieves an mIoU of 44.4% (single-scale) and 45.4% (multi-scale), attaining performance comparable to full-precision Swin-T and ConvNeXt-T. Even under aggressive 1-bit quantization, ViM-VQ maintains an mIoU above 40%, demonstrating its robustness in extremely low-bit settings.
@@ 5.3 Object Detection and Instance Segmentation
For object detection and instance segmentation on the MSCOCO dataset, ViM-VQ also achieves excellent results, as detailed in Table 3. At 2-bit quantization, ViM-VQ consistently outperforms full-precision Swin-T and ConvNeXt-T, achieving 46.0% box AP and 41.8% mask AP. Even under the extreme condition of 1-bit quantization, ViM-VQ effectively mitigates performance degradation, still obtaining a high 40.9% box AP and 35.8% mask AP.
@@
Table 5: Calibration efficiency of 2-bit quantization of tiny-scale Vision Mamba (Vim-T) on ImageNet-1K. $\mathbf{C}$, $n$, $E$, ’Mem’ represent the codebook size, the number of candidate codewords, the final epoch, and the memory usage for calibration, respectively. ’Acc-C’ and ’Acc-I’ represent the Top-1 accuracy (%) during calibration and inference.
@@ 5.4 Ablation Study
Table 4 presents an ablation study on the contribution of each component within our proposed ViM-VQ framework. The basic VQ method suffers from substantial quantization errors due to weight outliers, leading to a significant accuracy drop to 33.18%. By introducing the differentiable convex combination optimization algorithm, the accuracy improves by 38.24 percentage points, surpassing the performance of both DKM and VQ4DiT. This result highlights that updating both the convex combinations and the convex hulls effectively facilitates the search for the optimal codewords. Furthermore, incorporating the incremental vector quantization strategy provides an additional 0.75 percentage point improvement, demonstrating that the truncation error induced during the confirmation of optimal codewords is effectively mitigated.
@@
We develop a customized CUDA kernel for efficient matrix–vector multiplication. In our implementation, the codebooks are stored in constant GPU memory, while the assignments are fully loaded into CUDA blocks, enabling each thread to decode its corresponding assignment in parallel. The input is sequentially multiplied by the decoded weights, with the results accumulated in registers. As shown in Table 6, our customized CUDA kernel reduces inference time from 357s to 328s, nearly matching the inference speed of the original, full-precision ViM.
@@ 6 Conclusion
In this work, we identify and address key challenges encountered when applying vector quantization (VQ) to Visual Mamba networks (ViMs). These challenges include severe quantization errors caused by significant weight outliers, as well as the high calibration costs, suboptimal performance, and truncation errors of existing VQ methods. To overcome these issues, we propose ViM-VQ, a post-training VQ method specifically tailored for ViMs. ViM-VQ incorporates a fast convex combination optimization algorithm for efficient optimal codeword search and an incremental vector quantization strategy to confirm optimal codewords while mitigating truncation errors. Experimental results demonstrate that ViM-VQ significantly outperforms existing approaches across diverse visual tasks, offering a practical solution for deploying ViMs on edge devices.
7 Acknowledgements
This work was supported by the National Natural Science Foundation of China (Grant No. 62274142), Sino-German Mobility Programme (Grant No. M-0499), and Zhejiang University - Vivo Information Technology Joint Research Center.